# Insights into DeepSeek-V3: Scaling Challenges and Reflections on Hardware for AI Architectures

### v1

@@ 2.4. Technique Validation Methodology
Each acceleration technique undergoes rigorous empirical validation to evaluate its accuracy impact, including MLA, FP8 mixed-precision computation, and network co-designed MoE gate routing. Given the prohibitive cost of exhaustive ablation on full-scale models, we adopt a hierarchical and resource-efficient validation pipeline. Each technique is first validated extensively on small-scale models, followed by minimal large-scale tuning, and finally integrated in a single, comprehensive training run.
For instance, we first conducted fine-grained FP8 training ablation studies on both 16B and 230B DeepSeek-V2 models before final integration. Under these controlled settings, the relative accuracy loss compared to BF16 remains below 0.25%, attributable to our use of high-precision accumulation and fine-grained quantization strategies.
@@ 3.1. FP8 Mix-Precision Training
Quantization techniques such as GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2024) have been widely used to reduce bit-widths to 8-bit, 4-bit, or even lower, significantly reducing memory requirements. However, these techniques are primarily applied during inference to save memory, rather than in the training phase. NVIDIA’s Transformer Engine has supported FP8 mixed-precision training for some time, but prior to DeepSeek-V3, there were no open-source large models leveraging FP8 for training. Through deep collaboration between our infrastructure and algorithm teams, and after extensive experimentation and innovation, we developed an FP8-compatible training framework for MoE models. Figure 1 shows the computational components where FP8-precision forward and backward processes are utilized in the training pipeline. Fine-grained quantization is applied, i.e., tile-wise 1x128 quantization for activations and block-wise 128x128 quantization for model weights. Further technical details of our FP8 framework are documented in the DeepSeek-V3 technical report (DeepSeek-AI, 2024d), and our fine-grained FP8 GEMM implementation has been open-sourced in DeepGEMM (Zhao et al., 2025a).
@@ 3.1.1. Limitations:
  • FP8 Accumulation Precision: FP8 uses constrained accumulation precision in Tensor Cores, affecting the stability for training large models, particularly on NVIDIA Hopper GPUs. After aligning 32 mantissa products by right-shifting based on the maximum exponent, the Tensor Core only maintains their highest 13 fraction bits for addition, and truncates bits exceeding this range. Addition results are accumulated to FP22 registers (1 sign bit, 8 exponent bits, and 13 mantissa bits).
@@ 3.1.2. Suggestions:
  • Native Support for Fine-Grained Quantization: Hardware should natively support fine-grained quantization, enabling Tensor Cores to receive scaling factors and implement matrix multiplication with group scaling. In this way, the whole partial sum accumulation and dequantization can be completed directly inside Tensor Cores until the final result is produced, avoiding frequent data movements to reduce dequantization overhead. A notable industrial implementation of this approach is NVIDIA Blackwell’s support for microscaling data format (Rouhani et al., 2023), which exemplifies the practical benefits of native quantization at scale.
@@ 3.2. LogFMT: Communication Compression
By locally calculating the $min$ and $Step$, this data type supports dynamic representation range for different blocks, covering larger ranges or providing more precision, compared to static floating point formats. Besides, we find it is important to round in the original Linear space, instead of the Log space, for the unbiased activation quantization. We also constrain the $min$ to be larger than $max-log(2^{32})$, which means that the max representation range is similar to E5, a floating point with 5 exponents. We validate our LogFMT-nBit on dense language models with around 7 billion parameters, by quantifying the output of the residual branch to simulate the combine stage in MoE models. When setting $n=8$, sharing the same bits with FP8, the LogFMT-8Bit shows superior training accuracy compared to E4M3 or E5M2. After increasing the $n$ to 10 bits, we find it’s similar to the BF16 combine stage.
3.2.1. Limitations:
The initial purpose of using LogFMT is to apply it to activations during transmission or near activation functions, as it offers higher precision than FP8 with the same bit width. However, subsequent computations require reconversion to BF16 or FP8 to accommodate the Hopper GPU tensor cores’ data type. Due to insufficient GPU bandwidth for log/exp operations and excessive register pressure during encode/decode, if encode/decode operations are fused with all-to-all communication, the overhead can be substantial (50$\sim$100%). Therefore, although experimental results validate the effectiveness of this format, we do not employ it eventually.
@@ 4.2. Hardware-Aware Parallelism
  • Avoidance of Tensor Parallelism (TP): Tensor Parallelism is avoided during training due to its inefficiency under limited NVLink bandwidth. However, during inference, TP can still be selectively used to reduce latency and improve TPOT performance.
  • Enhanced Pipeline Parallelism (PP): DualPipe (DeepSeek-AI, 2025b) is employed to overlap attention and MoE computation with MoE communication. This also reduces pipeline bubbles and balances memory usage across GPUs, improving overall throughput. Additional details are available in the technical report (DeepSeek-AI, 2024d).
  • Accelerated Expert Parallelism (EP): With eight 400Gbps InfiniBand (IB) NICs, the system achieves all-to-all communication at speeds exceeding 40GB/s. Notably, our all-to-all EP implementation, DeepEP (Zhao et al., 2025b), is open-sourced, enabling highly efficient expert parallelism as discussed in the following subsection.
@@ 4.3. Model Co-Design: Node-Limited Routing
Consider a setup with 8 nodes (64 GPUs in total) and 256 routed experts (4 experts per GPU). For DeepSeek-V3, each token is routed to one shared expert and 8 routed experts. If its 8 target experts are distributed across all 8 nodes, the communication time over IB would be $8t$, where $t$ represents the time to send one token over IB. However, by leveraging the higher NVLink bandwidth, tokens routed to the same node can be sent once over IB and then forwarded via NVLink to other intra-node GPUs. The NVLink forwarding enables deduplication of the IB traffic. When the target experts for a given token are distributed across $M$ nodes, the deduplicated IB communication cost will be reduced to $Mt$ ($M<8$).
@@ 4.4.2. Suggestions:
We also recognize emerging interconnect protocols such as the Ultra Ethernet Consortium (UEC) (Consortium, 2023, 2024), Ultra Accelerator Link (UALink) (CONSORTIUM, 2025), both of which are poised to drive advancements in scale-up and scale-out communication. More recently, Unified Bus (UB) (Liao et al., 2025) has introduced a novel approach to scale-up and scale-out convergence. Section 6 further explores several technical innovations proposed by UEC and UALink. However, in this section, our primary focus is on achieving scale-up and scale-out convergence at the programming framework level.:
@@
    Dedicated Communication Co-Processor: Introduce a dedicated co-processor or programmable component—such as an I/O die—for handling network traffic. This component would offload packet processing from GPU SMs, preventing performance degradation. Besides, it should include hardware-accelerated memory copy capabilities for efficient buffer management.
@@ 5.1. Network Co-Design: Multi-Plane Fat-Tree
During the training of DeepSeek-V3, we deployed a Multi-Plane Fat-Tree (MPFT) scale-out network, as shown in Figure 3. Each node is equipped with eight GPUs and eight IB NICs, with each GPU–NIC pair assigned to a distinct network plane. Additionally, each node has a 400 Gbps Ethernet RoCE NIC connected to a separate storage network plane for accessing the 3FS (DeepSeek-AI, 2025c) distributed file system. In the scale-out network, we used 64-port 400G IB switches, enabling the topology theoretically supports up to 16,384 GPUs while retaining the cost and latency advantages of a two-layer network. However, due to policy and regulatory constraints, just over two thousand GPUs were ultimately deployed.
Figure 3. Eight-plane two-layer fat-tree scalue-out network: Each GPU and IB NIC pair belongs to one network plane. Cross-plane traffic must use another NIC and PCIe or NVLink for intra-node forwarding.
@@ 5.1.1. Advantages of Multi-Plane Fat-Tree Network
  • Subset of Multi-Rail Fat-Tree (MRFT): The MPFT topology constitutes a specific subset of the broader MRFT architecture. As a result, existing optimizations developed by NVIDIA and NCCL for Multi-Rail networks can be seamlessly leveraged within Multi-Plane network deployments. Furthermore, NCCL’s support for PXN (Mandakolathur and Jeaugey, 2022) technology addresses the inherent challenge of inter-plane isolation, enabling efficient communication even when direct interconnectivity between planes is absent.
  • Cost Efficiency: As shown in Table 3, the multi-plane network enables over 10k endpoints using a two-layer fat-tree (FT2) topology, significantly reducing network costs compared to a three-layer fat tree (FT3). The cost per endpoint is even slightly more competitive than the cost-efficient Slim Fly (SF) topology (Blach et al., 2025).
@@
Table 3. Network topology comparison. Cost estimates are derived from the methodology in the Slim Fly (SF) paper (Blach et al., 2025). DF denotes the canonical dragonfly topology (De Sensi et al., 2020; Kim et al., 2008; Rahman et al., 2019).
@@ 5.1.2. Performance Anlaysis
1. All-to-All Communication and EP Scenarios: As illustrated in Figure 5, the all-to-all performance of the multi-plane network is very similar to that of the single-plane multi-rail network. This performance parity can be attributed to NCCL’s PXN (Mandakolathur and Jeaugey, 2022) mechanism, which optimizes traffic forwarding via NVLink in multi-rail topologies. The multi-plane topology also benefits from this mechanism. As shown in Figure 6, the results of all-to-all communication tests conducted on 16 GPUs reveal negligible differences in latency between the MPFT and MRFT topologies.
@@
2. Training Throughput for DeepSeek-V3 Model: We also compare the training metrics of the DeepSeek-V3 model between MPFT and MRFT in Table 4. MFU (Model Flops Utilization) is calculated based on BF16 peak performance. Causal MFU only takes into account the flops of the lower triangle of the attention matrix (in line with FlashAttention(Dao et al., 2022; Dao, 2023)), while non-causal MFU includes the flops of the whole attention matrix (in line with Megatron (Korthikanti et al., 2023)). 1F, 1B, and 1W denote forward time, input backward time, and weight backward time, respectively. When training the V3 model on 2048 GPUs, the performance of MPFT is nearly identical to that of MRFT, with observed differences falling within normal fluctuations and measurement error.
@@
5.2. Low Latency Networks
In our model inference, large-scale EP relies heavily on all-to-all communication, which is highly sensitive to both bandwidth and latency. Consider a typical scenario discussed in Section 2.3.2, with a network bandwidth of 50GB/s, the data transfer should ideally take approximately $120\leavevmode\nobreak\ \mu\mathrm{s}$ Therefore, the intrinsic network latencies on the order of microseconds can critically impact system performance, making their effects non-negligible.
5.2.1. IB or RoCE
5.2.2. Recommendations for RoCE Improvements
  1. (1)
    Specialized Low-Latency RoCE Switches: We recommend that Ethernet vendors develop RoCE switches specifically optimized for RDMA workloads by removing unnecessary Ethernet features. The Slingshot architecture (De Sensi et al., 2020) exemplifies how Ethernet-based designs can achieve latency performance comparable to IB. Similarly, recent innovations from Broadcom (Broadcom, 2025), including the AI Forwarding Header (AIFH) and upcoming low-latency Ethernet switches, demonstrate the feasibility of high-performance Ethernet fabrics tailored for AI. We are looking forward to continuing innovation in this direction.
  2. (2)
    Optimized Route Policy: As shown in Figure 8, the default Equal-Cost Multi-Path (ECMP) routing policy in RoCE struggles to distribute traffic efficiently across interconnects, leading to severe congestion performance degradation in NCCL collective communication tests. LLM training traffic, such as in DP (Data Parallelism), tends to lack randomness, causing multiple flows to converge on the same interconnect link. In contrast, Adaptive Routing (AR) (Geoffray and Hoefler, 2008) can significantly enhance network performance by dynamically spraying packets across multiple paths. While static routing—based on manually configured route tables—can avoid link conflicts for specific destinations, it lacks flexibility. For large-scale all-to-all communication, adaptive routing offers superior performance and scalability.
  3. (3)
@@ 5.2.3. InfiniBand GPUDirect Async (IBGDA)
We utilize IBGDA (NVIDIA, 2022; Agostini et al., 2018) to reduce latency in network communications. Traditionally, network communication involves the creation of a CPU proxy thread: once the GPU has prepared the data, it must notify the CPU proxy, which then populates the control information for the work request (WR) and signals the NIC via a doorbell mechanism to initiate data transmission. This process introduces additional communication overhead.
IBGDA addresses this issue by allowing the GPU to directly fill the WR content and write to the RDMA doorbell MMIO address. By managing the entire control plane within the GPU, IBGDA eliminates the significant latency overhead associated with GPU-CPU communication. Moreover, when sending a large number of small packets, the control plane processor can easily become a bottleneck. Since GPUs have multiple parallel threads, the sender can leverage these threads to distribute the workload, thereby avoiding such bottlenecks. A range of works—including our DeepEP (Zhao et al., 2025b)—have leveraged IBGDA and reported substantial performance gains (Agostini et al., 2017; Chen et al., 2025; Zheng et al., 2025). We therefore advocate for such capabilities to be widely supported across accelerator devices.

### v2

@@ 2.4. Technique Validation Methodology
Each acceleration technique undergoes rigorous empirical validation to evaluate its accuracy impact, including MLA, FP8 mixed-precision computation, and network co-designed MoE gate routing. Given the prohibitive cost of exhaustive ablation on full-scale models, we adopt a hierarchical and resource-efficient validation pipeline. Each technique is first validated extensively on small-scale models, followed by minimal large-scale tuning, and finally integrated in a single, comprehensive training run. For instance, we first conducted fine-grained FP8 training ablation studies on both 16B and 230B DeepSeek-V2 models before final integration. Under these controlled settings, the relative accuracy loss compared to BF16 remains below 0.25%, attributable to our use of high-precision accumulation and fine-grained quantization strategies.
@@ 3.1. FP8 Mix-Precision Training
Quantization techniques such as GPTQ (frantar2022gptq) and AWQ (lin2023awq) have been widely used to reduce bit-widths to 8-bit, 4-bit, or even lower, significantly reducing memory requirements. However, these techniques are primarily applied during inference to save memory, rather than in the training phase. NVIDIA’s Transformer Engine has supported FP8 mixed-precision training for some time, but prior to DeepSeek-V3, there were no open-source large models leveraging FP8 for training. Through deep collaboration between our infrastructure and algorithm teams, and after extensive experimentation and innovation, we developed an FP8-compatible training framework for MoE models. Figure 1 shows the computational components where FP8-precision forward and backward processes are utilized in the training pipeline. Fine-grained quantization is applied, i.e., tile-wise 1x128 quantization for activations and block-wise 128x128 quantization for model weights. Further technical details of our FP8 framework are documented in the DeepSeek-V3 technical report (dsviii), and our fine-grained FP8 GEMM implementation has been open-sourced in DeepGEMM (deepgemm2025).
@@ 3.1.1. Limitations:
  • FP8 Accumulation Precision: FP8 uses constrained accumulation precision in Tensor Cores, affecting the stability for training large models, particularly on NVIDIA Hopper GPUs. After aligning 32 mantissa products by right-shifting based on the maximum exponent, the Tensor Core only maintains their highest 13 fraction bits for addition, and truncates bits exceeding this range. Addition results are accumulated to FP22 registers (1 sign bit, 8 exponent bits, and 13 mantissa bits)(zhang2025sageattention2efficientattentionthorough). The term “FP22” follows the naming used in the cited work, and the FP8 precision issue on Hopper GPUs had also been observed across industry by early 2024.
@@ 3.1.2. Suggestions:
  • Native Support for Fine-Grained Quantization: Hardware should natively support fine-grained quantization, enabling Tensor Cores to receive scaling factors and implement matrix multiplication with group scaling. In this way, the whole partial sum accumulation and dequantization can be completed directly inside Tensor Cores until the final result is produced, avoiding frequent data movements to reduce dequantization overhead. A notable industrial implementation of this approach is NVIDIA Blackwell’s support for microscaling data format (rouhani2023microscalingdataformatsdeep), which exemplifies the practical benefits of native quantization at scale.
@@ 3.2. LogFMT: Communication Compression
Besides these traditional floating point formats, we also tried a new data type, named Logarithmic Floating-Point Formats (LogFMT-nBit), where $n$ is the number of bits with the leading 1 bit as the sign bit $S$. By mapping the activations from the original Linear space to the Log space, the distribution of the activations is more uniform. To be specific, given a tile of elements, $x_{1},\cdots,x_{m}]$, which is 1x128 in our implementation, we take the absolute values and compute the logarithm of all the elements, and find the minimum $min=log(abs(x_{i}))$ and maximum $max=log(abs(x_{j}))$. The minimum is encoded as $S.00\cdots 01$ and the maximum is encoded as $S.11\cdots 11$, with an interval representing $Step=\frac{max-min}{2^{n-1}-2}$. Zero values are represented by $S.00\cdots 00$, specially. The left values are rounded to the nearest integer $K$ multiples of $Step$. The decoding process is simple by combining the sign bit and $exp^{min+Step\times(K-1)}$.
By locally calculating the minmin and $Step$, this data type supports dynamic representation range for different blocks, covering larger ranges or providing more precision, compared to static floating point formats. Besides, we find it is important to round in the original Linear space, instead of the Log space, for the unbiased activation quantization. We also constrain the minmin to be larger than $max-log(2^{32})$, which means that the max representation range is similar to E5, a floating point with 5 exponents. We validate our LogFMT-nBit on dense language models with around 7 billion parameters, by quantifying the output of the residual branch to simulate the combine stage in MoE models. When setting $n=8$, sharing the same bits with FP8, the LogFMT-8Bit shows superior training accuracy compared to E4M3 or E5M2. After increasing the $n$ to 10 bits, we find it’s similar to the BF16 combine stage.
3.2.1. Limitations:
@@ 4.2. Hardware-Aware Parallelism
  • Avoidance of Tensor Parallelism (TP): Tensor Parallelism is avoided during training due to its inefficiency under limited NVLink bandwidth. However, during inference, TP can still be selectively used to improve TTFT and TPOT performance.
  • Enhanced Pipeline Parallelism (PP): DualPipe (dualpipe) is employed to overlap attention and MoE computation with MoE communication. This also reduces pipeline bubbles and balances memory usage across GPUs, improving overall throughput. Additional details are available in the technical report (dsviii).
  • Accelerated Expert Parallelism (EP): With eight 400Gbps InfiniBand (IB) NICs, the system achieves all-to-all communication at speeds exceeding 40GB/s. Notably, our all-to-all EP implementation, DeepEP (deepep2025), is open-sourced, enabling highly efficient expert parallelism as discussed in the following subsection.
Figure 2. H800 node interconnection.
@@ 4.3. Model Co-Design: Node-Limited Routing
Consider a setup with 8 nodes (64 GPUs in total) and 256 routed experts (4 experts per GPU). For DeepSeek-V3, each token is routed to one shared expert and 8 routed experts. If its 8 target experts are distributed across all 8 nodes, the communication time over IB would be $8t$, where $t$ represents the time to send one token over IB. However, by leveraging the higher NVLink bandwidth, tokens routed to the same node can be sent once over IB and then forwarded via NVLink to other intra-node GPUs. The NVLink forwarding enables deduplication of the IB traffic. When the target experts for a given token are distributed across $M$ nodes, the deduplicated IB communication cost will be reduced to MtMt ($M<8$).
@@ 4.4.2. Suggestions:
We also recognize emerging interconnect protocols such as the Ultra Ethernet Consortium (UEC) (uec_overview; uec_update1), Ultra Accelerator Link (UALink) (ualink_white_paper), both of which are poised to drive advancements in scale-up and scale-out communication. More recently, Unified Bus (UB) (liao2025ubmeshhierarchicallylocalizedndfullmesh) has introduced a novel approach to scale-up and scale-out convergence. Section 6 further explores several technical innovations proposed by UEC and UALink. However, in this section, our primary focus is on achieving scale-up and scale-out convergence at the programming framework level:
@@
    Dedicated Communication Co-Processor: Introduce a dedicated co-processor or programmable component—such as an I/O die—for handling network traffic. This component should offload packet processing from GPU SMs, provide hardware-accelerated memory copy for efficient buffer management, and, crucially, accelerate memory load/store operations in a manner similar to TMA (Tensor Memory Accelerator), thereby saturating bandwidth with minimal resource consumption.
@@ 4.5.1. Limitations:
Figure 3. Eight-plane two-layer fat-tree scale-out network: Each GPU and IB NIC pair belongs to one network plane. Cross-plane traffic must use another NIC and PCIe or NVLink for intra-node forwarding.
@@ 5.1. Network Co-Design: Multi-Plane Fat-Tree
During the training of DeepSeek-V3, we deployed a Multi-Plane Fat-Tree (MPFT) scale-out network, as shown in Figure 3. Each node is equipped with eight GPUs and eight IB NICs, with each GPU–NIC pair assigned to a distinct network plane. Additionally, each node has a 400 Gbps Ethernet RoCE NIC connected to a separate storage network plane for accessing the 3FS (ds3fs) distributed file system. In the scale-out network, we used 64-port 400G IB switches, enabling the topology theoretically supports up to 16,384 GPUs while retaining the cost and latency advantages of a two-layer network. However, due to policy and regulatory constraints, just over two thousand GPUs were ultimately deployed.
@@ 5.1.1. Advantages of Multi-Plane Fat-Tree Network
  • Subset of Multi-Rail Fat-Tree (MRFT): The MPFT topology constitutes a specific subset of the broader MRFT architecture. As a result, existing optimizations developed by NVIDIA and NCCL for Multi-Rail networks can be seamlessly leveraged within Multi-Plane network deployments. Furthermore, NCCL’s support for PXN (nccl_pxn) technology addresses the inherent challenge of inter-plane isolation, enabling efficient communication even when direct interconnectivity between planes is absent.
  • Cost Efficiency: As shown in Table 3, the multi-plane network enables over 10k endpoints using a two-layer fat-tree (FT2) topology, significantly reducing network costs compared to a three-layer fat tree (FT3). The cost per endpoint is even slightly more competitive than the cost-efficient Slim Fly (SF) topology (10.5555/3691825.3691882).
@@
Table 3. Network topology comparison. Cost estimates are derived from the methodology in the Slim Fly (SF) paper (10.5555/3691825.3691882). DF denotes the canonical dragonfly topology (9355230; 4556717; 10.1145/3295500.3356208).
@@ 5.1.2. Performance Analysis
1. All-to-All Communication and EP Scenarios: As illustrated in Figure 5, the all-to-all performance of the multi-plane network is very similar to that of the single-plane multi-rail network. This performance parity can be attributed to NCCL’s PXN (nccl_pxn) mechanism, which optimizes traffic forwarding via NVLink in multi-rail topologies. The multi-plane topology also benefits from this mechanism. As shown in Figure 6, the results of all-to-all communication tests conducted on 16 GPUs reveal negligible differences in latency between the MPFT and MRFT topologies.
@@
2. Training Throughput for DeepSeek-V3 Model: We also compare the training metrics of the DeepSeek-V3 model between MPFT and MRFT in Table 4. MFU (Model Flops Utilization) is calculated based on BF16 peak performance. Causal MFU only takes into account the flops of the lower triangle of the attention matrix (in line with FlashAttention(dao2022flashattention; dao2023flashattention2)), while non-causal MFU includes the flops of the whole attention matrix (in line with Megatron (megatron3)). 1F, 1B, and 1W denote forward time, input backward time, and weight backward time, respectively. When training the V3 model on 2048 GPUs, the performance of MPFT is nearly identical to that of MRFT, with observed differences falling within normal fluctuations and measurement error.
5.2. Low Latency Networks
In our model inference, large-scale EP relies heavily on all-to-all communication, which is highly sensitive to both bandwidth and latency. Consider a typical scenario discussed in Section 2.3.2, with a network bandwidth of 50GB/s, the data transfer should ideally take approximately $120\penalty 10000\ \mu\mathrm{s}$ . Therefore, the intrinsic network latencies on the order of microseconds can critically impact system performance, making their effects non-negligible.
5.2.1. IB or RoCE
@@
5.2.2. Recommendations for RoCE Improvements
  1. (1)
    Specialized Low-Latency RoCE Switches: We recommend that Ethernet vendors develop RoCE switches specifically optimized for RDMA workloads by removing unnecessary Ethernet features. The Slingshot architecture (9355230) exemplifies how Ethernet-based designs can achieve latency performance comparable to IB. Similarly, recent innovations from Broadcom (brcm_eth_scale), including the AI Forwarding Header (AIFH) and upcoming low-latency Ethernet switches, demonstrate the feasibility of high-performance Ethernet fabrics tailored for AI. We are looking forward to continuing innovation in this direction.
  2. (2)
    Optimized Route Policy: As shown in Figure 8, the default Equal-Cost Multi-Path (ECMP) routing policy in RoCE struggles to distribute traffic efficiently across interconnects, leading to severe congestion performance degradation in NCCL collective communication tests. LLM training traffic, such as in DP (Data Parallelism), tends to lack randomness, causing multiple flows to converge on the same interconnect link. In contrast, Adaptive Routing (AR) (4618589) can significantly enhance network performance by dynamically spraying packets across multiple paths. While static routing—based on manually configured route tables—can avoid link conflicts for specific destinations, it lacks flexibility. For large-scale all-to-all communication, adaptive routing offers superior performance and scalability.
  3. (3)
@@ 5.2.3. InfiniBand GPUDirect Async (IBGDA)
We utilize IBGDA (nvshmem_ibgda; AGOSTINI201828) to reduce latency in network communications. Traditionally, network communication involves the creation of a CPU proxy thread: once the GPU has prepared the data, it must notify the CPU proxy, which then populates the control information for the work request (WR) and signals the NIC via a doorbell mechanism to initiate data transmission. This process introduces additional communication overhead.
IBGDA addresses this issue by allowing the GPU to directly fill the WR content and write to the RDMA doorbell MMIO address. By managing the entire control plane within the GPU, IBGDA eliminates the significant latency overhead associated with GPU-CPU communication. Moreover, when sending a large number of small packets, the control plane processor can easily become a bottleneck. Since GPUs have multiple parallel threads, the sender can leverage these threads to distribute the workload, thereby avoiding such bottlenecks. A range of works—including our DeepEP (deepep2025)—have leveraged IBGDA and reported substantial performance gains (7973709; chen2025efficientheterogeneouslargelanguage; zheng2025tilelinkgeneratingefficientcomputecommunication). We therefore advocate for such capabilities to be widely supported across accelerator devices.